The pOH of pure water at 40oC is 6.8. What is the hydronium concentration, [H3O+], in pure water at this temperature?
Answer #1

use:

pOH = -log [OH-]

6.8 = -log [OH-]

[OH-] = 1.585*10^-7 M

use:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/1.585*10^-7

[H3O+] = 6.31*10^-8 M

Answer: 6.31*10^-8 M